OpenStudy (ksaimouli):

solve the inequalities 2x-1/3x-2<=0

OpenStudy (ksaimouli):



OpenStudy (lgbasallote):

start by cross multiplying \[\huge \implies 2x - 1 \le 0 \times (3x - 1)\] simplify... \[\huge \implies 2x - 1 \le 0\] add 1 to both sides \[\huge \implies 2x \le 1\] divide both sides by 2 \[\huge \implies x \le \frac 12\] doe that help?

OpenStudy (ksaimouli):

yes thx

OpenStudy (ksaimouli):

what about x^2+2x-3<=0